You own a bond that has a duration of 6 years. Interest rates are currently 7%, but you believe the Fed is about to increase interest rates by 22 basis points. Your predicted price change on this bond is ________.
Answer #1

Solution:

D* = Given duration /(1+Interest rate) = 6 / (1+7%)

D* = 6/1.07

D* = 5.607476636

----

Now,

Percentage change in price = -D* x Change in yield

Percentage change in price = -5.607476636 x 0.22%

Percentage change in price = -1.2336%

Percentage change in price = -1.23%

Answer in two decimals: -1.23%

------

Hence,

You own a bond that has a duration of 6 years. Interest rates are currently 7%, but you believe the Fed is about to increase interest rates by 22 basis points. Your predicted price change on this bond is -1.23%